# The barrier Ramsey theorem

### v2

@@ 2.1. Systems of Fundamental Sequences
Let $\Gamma$ be an ordinal. If we fix a fundamental sequence for each ordinal $\alpha<\Gamma$ we speak of a system of fundamental sequences on $\Gamma$.
@@
A system of fundamental sequences on $\Gamma$ is regular if for each ordinal $\beta_{0}+\omega^{\beta_{1}}<\Gamma$ with $\beta_{0}\mathrel{\geq\!\!\!\geq}\omega^{\beta_{1}}$ (i.e. $\beta_{0}$ are the leading terms in the Cantor normal form, which has another term, of $\beta_{0}+\omega^{\beta_{1}}$) and each $n$ it holds that
@@
A system of fundamental sequences on $\Gamma$ is nested if it is never the case for $\gamma<\beta<\Gamma$ and $n>1$ that
@@
Regularity is a very natural property that is satisfied by most systems of fundamental sequences that have been studied. Nestedness is a weakening of the Bachmann property which states that for ordinals $\gamma<\beta<\Gamma$ it is never the case that $\gamma>\beta[n]>\gamma[1]$. The Bachmann property was introduced in [Bac67] and studied in [PTV91]. An example of a system of fundamental sequences that satisfies the Bachmann property (but, incidentally, is not regular) is provided in [FDW24].
For the rest of Subsection 2.1 we fix a system of fundamental sequences on some ordinal $\Gamma$. We also assume to work only with ordinals $\Gamma$.
@@
Given a system of fundamental sequences on some ordinal $\zeta$ (which we denote as $\xi\lceil n\rceil$ for $\xi<\zeta$), we define a system of fundamental sequences on $\Gamma_{\zeta}$ as follows.
@@
    If $0<\xi<\zeta$ then $\Gamma_{\xi}[n]=\varphi_{\varphi_{\ddots\varphi_{\Gamma_{\xi\lceil n\rceil}+1}(0)\iddots}(0)}(0)$, with $\varphi$ iterated $n+1$ times.
@@ 2.2. Fronts, Blocks and Barriers
A subfront (respectively, a subblock, a subbarrier) is a subset of a front which is still a front (respectively, a block, a barrier). If $\mathcal{B}$ is a block and $\mathcal{B}^{\prime}\subseteq\mathcal{B}$ is a subblock of $\mathcal{B}$ then it must be $\mathcal{B}^{\prime}=\{s\in\mathcal{B}:s\subseteq X\}$ for some infinite $X\subseteq\operatorname{base}(\mathcal{B})$. Vice versa, for each infinite $X\subseteq\operatorname{base}(\mathcal{B})$ we obtain a subblock of $\mathcal{B}$ by restricting to the elements of $\mathcal{B}$ which are subsets of $X$. It is straightforward to see that a smooth block is automatically a barrier. It is also immediate that any subblock of a (smooth) barrier is a (smooth) barrier.
@@
Now assume that $\mathcal{B}$ is smooth and fix $s,t\in T(\mathcal{B})$ of the same length and such that $0pt(s)<0pt(t)$. We proceed by induction on $0pt(s)$. If $0pt(s)=0$ then $0pt(t)>0$ and so $s\in\mathcal{B}$ and $t\in T(\mathcal{B})\setminus\mathcal{B}$. Let $u\in\mathcal{B}$ be an extension of $t$. Then $|s|=|t|<|u|$ and so by $1$ there exists $i<|s|$ such that $s(i)<u(i)=t(i)$. Suppose now that $0pt(s)=\alpha>0$. Let $n>\max(s\cup t)$ be such that $0pt(t{}^{\smallfrown}\langle n\rangle)\geq\alpha$. Then $0pt(s{}^{\smallfrown}\langle n\rangle)<\alpha\leq 0pt(t{}^{\smallfrown}\langle n\rangle)$. By inductive hypothesis there exists $i<|s{}^{\smallfrown}\langle n\rangle|$ such that $s{}^{\smallfrown}\langle n\rangle(i)<t{}^{\smallfrown}\langle n\rangle(i)$. Notice that it must be $i<|s|$ because $s{}^{\smallfrown}\langle n\rangle(|s|)=n=t{}^{\smallfrown}\langle n\rangle(|s|)$. Therefore $s(i)<t(i)$. ∎
@@
Let $X\subseteq\operatorname{base}(\mathcal{B})$ be infinite and let $s\in\mathcal{B}$ be such that $s\sqsubset X$. Let $u\sqsubseteq s$ be longest such that $0pt_{\mathcal{B}}(u)\leq\beta_{0}$. Notice that $u\neq\langle\rangle$ and hence $u\in\overline{\mathcal{B}}_{1}$. Thus $\overline{\mathcal{B}}_{1}$ is a block.
@@
We are left to prove that each element of $\overline{\mathcal{B}}_{0}\oplus\overline{\mathcal{B}}_{1}$ extends some element of $\mathcal{B}$. Let $t{}^{\smallfrown}s\in\overline{\mathcal{B}}_{0}\oplus\overline{\mathcal{B}}_{1}$ with $t\in\overline{\mathcal{B}}_{1}$ and $s\in\overline{\mathcal{B}}_{0}$. Towards a contradiction, suppose that $t{}^{\smallfrown}s\in T(\mathcal{B})\setminus\mathcal{B}$ and let $s^{\prime}$ be such that $s\sqsubset s^{\prime}$ and $t{}^{\smallfrown}s^{\prime}\in\mathcal{B}$. Then $s^{\prime}\in\mathcal{B}_{t}$ and so some extension of $s^{\prime}$ belongs to $\mathcal{B}_{0}$. Since we are assuming $s\in\overline{\mathcal{B}}_{0}$, this contradicts that $\mathcal{B}_{0}$ is a block. ∎
@@ 3. Bridges Between Systems and Barriers
For the rest of this section, fix a system of fundamental sequences on some ordinal $\Gamma$ and an infinite set $M\subseteq\mathbb{N}$ with $\min M>1$222the restriction to natural numbers larger than $1$ mirrors our definition of nestedness.. If $\alpha<\Gamma$, let $\mathcal{B}^{\alpha}=[M]^{\alpha}$ be the set of $\alpha$-size subsets of $M$.
@@
For each $\alpha<\Gamma$, $\mathcal{B}^{\alpha}$ is a front. Moreover, if the system is nested every $\mathcal{B}^{\alpha}$ is a smooth barrier.
@@
If the sequence is non decreasing there exist $n<m$ with $\alpha[m]<\alpha[n]$, i.e. $0pt(\mathcal{B}_{m})<0pt(\mathcal{B}_{n})$. Let $\mathcal{B}_{n}^{\prime}$ be the restriction of $\mathcal{B}_{n}$ to $\operatorname{base}(\mathcal{B}_{m})$, which is a smooth barrier with the same height as $\mathcal{B}_{n}$ by Lemma 2.23. By Lemma 2.24 there exist $s\in\mathcal{B}_{m}$ and $t\in\mathcal{B}_{n}^{\prime}$ such that $s\sqsubset t$. The sequences $\langle m\rangle{}^{\smallfrown}s$ and $\langle n\rangle{}^{\smallfrown}t$ belong to $\mathcal{B}$ and contradict its smoothness.
@@ 4. Ramsey Theorem for Barriers
Let $k\in\mathbb{N}$ and let $\mathcal{A}_{0},\ldots,\mathcal{A}_{k-1},\mathcal{C}$ be fronts such that each of $\operatorname{base}(\mathcal{A}_{i})$ is a final segment of $\operatorname{base}(\mathcal{C})$. Then every infinite $Y\subseteq\bigcap_{i<k}\operatorname{base}(\mathcal{A}_{i})$ has an initial segment $s_{Y}$ such that for each coloring $c\colon[s_{Y}]^{\mathcal{C}}\rightarrow k$ there exists a homogeneous $(\mathcal{C}\oplus\mathcal{A}_{i})$-size subset of $s_{Y}$ of color $i$ for some $i<k$.
@@
Notice that in the previous lemma, if at least one between $\mathcal{A}$ and $\mathcal{C}$ is a block, then $s_{Y}$ cannot be $\langle\rangle$.
We can now obtain what we call the barrier Ramsey theorem444beware that the terminology barrier Ramsey theorem is used in [CGHP24] for a different statement., which allows us to introduce the Ramsey ordinals in our framework.
@@
For any countable ordinals $\alpha$ and $\gamma$ such that $\gamma<\alpha$ and $\alpha\geq\omega$ we have $\operatorname{Ram}(\alpha)^{1+\gamma}_{<\omega}=\varphi_{\log\gamma}(\alpha\cdot\omega)$.
@@ 6. The Lower Bound
We now show how Theorem 6.1 yields our lower bound. Fix $\gamma<\alpha<\Gamma_{\zeta}$ and $\nu<\varphi_{\log\gamma}(\alpha\mathbin{{\times}\!\!\!\!{\times}}\omega)<\Gamma_{\zeta}$ ordinals. Pick $k\in\mathbb{N}$ and $\mu$ an ordinal such that $\nu<\mu<\varphi_{\log\gamma}(\alpha\mathbin{{\times}\!\!\!\!{\times}}k)$. Apply Theorem 6.1 to $k,\gamma,\alpha,\mu$ to get an infinite set $M$ (we may assume $\min M>1$). Now let $\mathcal{A}=[M]^{\alpha}$, $\mathcal{C}=[M]^{1+\gamma}$ and $\mathcal{D}=[M]^{1\uplus\mu}$. By Theorem 3.6 $\mathcal{A}$, $\mathcal{C}$ and $\mathcal{D}$ are $\mathsf{SD}$-barrier. Let $\mathcal{B}$ be any $\mathsf{SD}$-barrier of height $\nu$ with $\operatorname{base}(\mathcal{B})=M$. Since $0pt(\mathcal{B})=\nu<1+\mu=0pt(\mathcal{D})$, by Lemma 2.24 there are $s\in\mathcal{B}$ and $t\in\mathcal{D}$ such that $s\sqsubset t$. Then $s{}^{\smallfrown}\langle t(|s|)\rangle$ inherits a bad coloring of its $(1+\gamma)$-size sets with $k+3$ colors from $t$ and from the fact that $t\in\mathcal{D}$. This means that $\mathcal{B}\nrightarrow(\mathcal{A})^{\mathcal{C}}_{k+3}$ and, since $\mathcal{B}$ is a generic $\mathsf{SD}$-barrier of height $\nu$ with $\operatorname{base}(\mathcal{B})=\operatorname{base}(\mathcal{A})$, we conclude.
@@
  • If $\beta=\omega^{\beta_{0}}+\ldots+\omega^{\beta_{n-1}}$ with $n>0$ and $\sigma_{0}\geq\ldots\geq\sigma_{n}$, let $\operatorname{Sub}(\beta)=\{\beta\}+\operatorname{Sub}(\omega^{\beta_{0}})+\ldots+\operatorname{Sub}(\omega^{\beta_{n-1}})$.
@@
$\bar{p}_{1}(\beta_{0},\beta_{1},\dots,\beta_{l})=(\overline{\beta_{0},\beta_{1}},\overline{\beta_{1},\beta_{2}},\ldots,\overline{\beta_{k},0}).$
@@
Notice that $0\in S(\tau)$ for every $\tau$, which takes care of the case when $\tau=A(0)\leq A(1)$. So we can assume $A(0)>A(1)$.
@@
When $\nu=\gamma$ this shows that the coloring $\overline{c}$ defined above can actually be regarded as a coloring of the $(1+\gamma)$-size subsets of $s$.
@@
By definition of $p_{\zeta}(t^{-0})$, we know that the Cantor normal forms of the ordinals $p_{\zeta-1}(t^{-0})$ and $p_{\zeta-1}(t^{-1})$ are equal up to a certain term $\omega^{\alpha_{n_{0}}}$, and then in $p_{\zeta-1}(t^{-0})$ we have the term $\omega^{p_{\zeta}(t^{-0})}$, while the tail $\tau_{1}$ of $p_{\zeta-1}(t^{-1})$ is strictly smaller than $\omega^{p_{\zeta}(t^{-0})}$. Since $p_{\zeta}(t^{-0})\leq p_{\zeta}(t^{-1})$, we must have that $p_{\zeta}(t^{-1})$ is an exponent of the Cantor normal form of $p_{\zeta-1}(t^{-1})$ that shows up before $n_{0}$. This implies that $p_{\zeta}(t^{-1})$ is one of the exponents of $p_{\zeta-1}(t^{-0})$ too. Since $\omega^{p_{\zeta}(t^{-0})}>\tau_{1}$, we must have that $\omega^{p_{\zeta}(t^{-1})}$ comes strictly before $\omega^{p_{\zeta}(t^{-0})}$ in the Cantor normal form of $p_{\zeta-1}(t^{-0})$. That is, even if they are equal as ordinals, they are different subterms of $\omega^{p_{\zeta}(t^{-0})}$. Continuing like this we get that each $p_{\zeta}(t^{-i})$ is a different exponent in the Cantor normal form of $p_{\zeta-1}(t^{-0})$}. Finally, recall that $p_{\zeta-1}(t^{-0})$} is already a subterm of $T(t(0))$ by Lemma 6.5. Therefore, we get that that each $p_{\zeta}(t^{-i})$ for $i\leq|t_{\alpha}|$ originates from a distinct occurrence of an element of $\operatorname{Sub}(T(t(0)))$, as required.

### v3

# arXiv:2505.02544v3[math.LO] 02 Mar 2026

@@ 2.1. Systems of Fundamental Sequences
Let $\Gamma$ be an ordinal. If we fix a fundamental sequence for each ordinal $\alpha\leq\Gamma$ we speak of a system of fundamental sequences on $\Gamma$.
@@
A system of fundamental sequences on $\Gamma$ is regular if for each ordinal $\beta_{0}+\omega^{\beta_{1}}\leq\Gamma$ with $\beta_{0}\mathrel{\geq\!\!\!\geq}\omega^{\beta_{1}}$ (i.e. $\beta_{0}$ are the leading terms in the Cantor normal form, which has another term, of $\beta_{0}+\omega^{\beta_{1}}$) and each $n$ it holds that
@@
A system of fundamental sequences on $\Gamma$ is nested if it is never the case for $\gamma<\beta\leq\Gamma$ and $n>1$ that
@@
Regularity is a very natural property that is satisfied by most systems of fundamental sequences that have been studied. Nestedness is a weakening of the Bachmann property which states that for ordinals $\gamma<\beta\leq\Gamma$ it is never the case that $\gamma>\beta[n]>\gamma[1]$. The Bachmann property was introduced in [Bac67] and studied in [PTV91]. An example of a system of fundamental sequences that satisfies the Bachmann property (but, incidentally, is not regular) is provided in [FDW24].
For the rest of Subsection 2.1 we fix a system of fundamental sequences on some ordinal $\Gamma$. We also assume to work only with ordinals $\leq\Gamma$.
@@
Given a system of fundamental sequences on some ordinal $\zeta$ (which we denote as $\xi\lceil n\rceil$ for $\xi\leq\zeta$), we define a system of fundamental sequences on $\Gamma_{\zeta}$ as follows.
@@
    If $0<\xi\leq\zeta$ then $\Gamma_{\xi}[n]=\varphi_{\varphi_{\ddots\varphi_{\Gamma_{\xi\lceil n\rceil}+1}(0)\iddots}(0)}(0)$, with $\varphi$ iterated $n+1$ times.
@@ 2.2. Fronts, Blocks and Barriers
A subfront (respectively, a subblock, a subbarrier) is a subset of a front which is still a front (respectively, a block, a barrier). If $\mathcal{B}$ is a block and $\mathcal{B}^{\prime}\subseteq\mathcal{B}$ is a subblock of $\mathcal{B}$ then it must be $\mathcal{B}^{\prime}=\{s\in\mathcal{B}:s\subseteq X\}$ for some infinite $X\subseteq\operatorname{base}(\mathcal{B})$. Vice versa, for each infinite $X\subseteq\operatorname{base}(\mathcal{B})$ we obtain a subblock of $\mathcal{B}$ by restricting to the elements of $\mathcal{B}$ which are subsets of $X$: we denote it by $\mathcal{B}\restriction X$. It is straightforward to see that a smooth block is automatically a barrier. It is also immediate that any subblock of a (smooth) barrier is a (smooth) barrier.
@@
Now assume that $\mathcal{B}$ is smooth and fix $s,t\in T(\mathcal{B})$ of the same length and such that $0pt(s)<0pt(t)$. We proceed by induction on $0pt(s)$. If $0pt(s)=0$ then $0pt(t)>0$ and so $s\in\mathcal{B}$ and $t\in T(\mathcal{B})\setminus\mathcal{B}$. Let $u\in\mathcal{B}$ be an extension of $t$. Then $|s|=|t|<|u|$ and so by 1. there exists $i<|s|$ such that $s(i)<u(i)=t(i)$. Suppose now that $0pt(s)=\alpha>0$. By Lemma 2.23 $0pt(\mathcal{B}_{t})=0pt(\mathcal{B}_{t}\restriction X)$ where $X=\operatorname{base}(\mathcal{B})\setminus\{0,\ldots,\max(s\cup t)\}$, hence there exists $n>\max(s\cup t)$ such that $0pt(t{}^{\smallfrown}\langle n\rangle)\geq\alpha$. Then $0pt(s{}^{\smallfrown}\langle n\rangle)<\alpha\leq 0pt(t{}^{\smallfrown}\langle n\rangle)$. By inductive hypothesis there exists $i<|s{}^{\smallfrown}\langle n\rangle|$ such that $s{}^{\smallfrown}\langle n\rangle(i)<t{}^{\smallfrown}\langle n\rangle(i)$. Notice that it must be $i<|s|$ because $s{}^{\smallfrown}\langle n\rangle(|s|)=n=t{}^{\smallfrown}\langle n\rangle(|s|)$. Therefore $s(i)<t(i)$. ∎
@@
Let $X\subseteq\operatorname{base}(\mathcal{B})$ be infinite and let $s\in\mathcal{B}$ be such that $s\sqsubset X$. Let $u\sqsubseteq s$ be shortest such that $0pt_{\mathcal{B}}(u)\leq\beta_{0}$. Notice that $u\neq\langle\rangle$ and hence $u\in\overline{\mathcal{B}}_{1}$. Thus $\overline{\mathcal{B}}_{1}$ is a block.
@@
We are left to prove that each element of $\overline{\mathcal{B}}_{0}\oplus\overline{\mathcal{B}}_{1}$ extends some element of $\mathcal{B}$. Let $t{}^{\smallfrown}s\in\overline{\mathcal{B}}_{0}\oplus\overline{\mathcal{B}}_{1}$ with $t\in\overline{\mathcal{B}}_{1}$ and $s\in\overline{\mathcal{B}}_{0}$. Towards a contradiction, suppose that $t{}^{\smallfrown}s\in T(\mathcal{B})\setminus\mathcal{B}$ and let $s^{\prime}$ be such that $s\sqsubset s^{\prime}$ and $t{}^{\smallfrown}s^{\prime}\in\mathcal{B}$. Then $s^{\prime}\in\mathcal{B}_{t}$ and so some extension of $s^{\prime}$ belongs to $\overline{\mathcal{B}}_{0}$. Since we are assuming $s\in\overline{\mathcal{B}}_{0}$, this contradicts that $\mathcal{B}_{0}$ is a block. ∎
@@
Remark 2.33.
If we restrict a $\mathsf{SD}$-barrier $\mathcal{B}$ to a final segment $X$ of its base, we still get a $\mathsf{SD}$-barrier. This is because smoothness is clearly preserved since we have a subbarrier, height is the same by Lemma 2.23 and any decomposition $\mathcal{B}_{0},\ldots,\mathcal{B}_{n}$ of $\mathcal{B}$ yields a decomposition $\mathcal{B}_{0}\restriction X,\ldots,\mathcal{B}_{n}\restriction X$ of $\mathcal{B}\restriction X$.
@@ 3. Bridges Between Systems and Barriers
For the rest of this section, fix a system of fundamental sequences on some ordinal $\Gamma$ and an infinite set $M\subseteq\mathbb{N}$ with $\min M>1$222the restriction to natural numbers larger than $1$ mirrors our definition of nestedness.. If $\alpha\leq\Gamma$, let $\mathcal{B}^{\alpha}=[M]^{\alpha}$ be the set of $\alpha$-size subsets of $M$.
@@
For each $\alpha\leq\Gamma$, $\mathcal{B}^{\alpha}$ is a front. Moreover, if the system is nested every $\mathcal{B}^{\alpha}$ is a smooth barrier.
@@
If the sequence is not non decreasing there exist $n<m$ with $\alpha[m]<\alpha[n]$, i.e. $0pt(\mathcal{B}_{m})<0pt(\mathcal{B}_{n})$. Let $\mathcal{B}_{n}^{\prime}$ be the restriction of $\mathcal{B}_{n}$ to $\operatorname{base}(\mathcal{B}_{m})$, which is a smooth barrier with the same height as $\mathcal{B}_{n}$ by Lemma 2.23. By Lemma 2.24 there exist $s\in\mathcal{B}_{m}$ and $t\in\mathcal{B}_{n}^{\prime}$ such that $s\sqsubset t$. The sequences $\langle m\rangle{}^{\smallfrown}s$ and $\langle n\rangle{}^{\smallfrown}t$ belong to $\mathcal{B}$ and contradict its smoothness.
@@ 4. Ramsey Theorem for Barriers
Let $k\in\mathbb{N}$ and let $\mathcal{A}_{0},\ldots,\mathcal{A}_{k-1},\mathcal{C}$ be fronts such that each of $\operatorname{base}(\mathcal{A}_{i})$ is a final segment of $\operatorname{base}(\mathcal{C})$. Then each infinite $Y\subseteq\bigcap_{i<k}\operatorname{base}(\mathcal{A}_{i})$ has an initial segment $s_{Y}$ such that for each coloring $c\colon[s_{Y}]^{\mathcal{C}}\rightarrow k$ there exists a homogeneous $(\mathcal{C}\oplus\mathcal{A}_{i})$-size subset of $s_{Y}$ of color $i$ for some $i<k$.
@@
Notice that in the previous lemma, if at least one between $\mathcal{A}_{i}$ and $\mathcal{C}$ is a block, then $s_{Y}$ cannot be $\langle\rangle$.
We can now obtain what we call the barrier Ramsey theorem444beware that the terminology barrier Ramsey theorem is used in [CGLHP] for a different statement., which allows us to introduce the Ramsey ordinals in our framework.
@@
For any countable ordinals $\alpha$ and $\gamma$ such that $\gamma<\alpha<\Gamma_{\zeta}$ and $\alpha\geq\omega$ we have $\operatorname{Ram}(\alpha)^{1+\gamma}_{<\omega}=\varphi_{\log\gamma}(\alpha\cdot\omega)$.
@@ 6. The Lower Bound
We now show how Theorem 6.1 yields our lower bound. Fix $\gamma<\alpha<\Gamma_{\zeta}$ and $\nu<\varphi_{\log\gamma}(\alpha\mathbin{{\times}\!\!\!\!{\times}}\omega)<\Gamma_{\zeta}$ ordinals. Pick $k\in\mathbb{N}$ and $\mu$ an ordinal such that $\nu<\mu<\varphi_{\log\gamma}(\alpha\mathbin{{\times}\!\!\!\!{\times}}k)$. Apply Theorem 6.1 to $k,\gamma,\alpha,\mu$ to get an infinite set $M$ (we may assume $\min M>1$). Now let $\mathcal{A}=[M]^{\alpha}$, $\mathcal{C}=[M]^{1+\gamma}$ and $\mathcal{D}=[M]^{1\uplus\mu}$. Notice that $\mathcal{A}$ and $\mathcal{C}$ do not depend on $\nu$. By Theorem 3.6 $\mathcal{A}$, $\mathcal{C}$ and $\mathcal{D}$ are $\mathsf{SD}$-barrier. Let $\mathcal{B}$ be any $\mathsf{SD}$-barrier of height $\nu$ with $\operatorname{base}(\mathcal{B})=M$. Since $0pt(\mathcal{B})=\nu<1+\mu=0pt(\mathcal{D})$, by Lemma 2.24 there are $s\in\mathcal{B}$ and $t\in\mathcal{D}$ such that $s\sqsubset t$. Then $s{}^{\smallfrown}\langle t(|s|)\rangle$ inherits a bad coloring of its $(1+\gamma)$-size sets with $k+3$ colors from $t$ and from the fact that $t\in\mathcal{D}$. This means that $\mathcal{B}\nrightarrow(\mathcal{A})^{\mathcal{C}}_{k+3}$ and, since $\mathcal{B}$ is a generic $\mathsf{SD}$-barrier of height $\nu$ with $\operatorname{base}(\mathcal{B})=\operatorname{base}(\mathcal{A})$, we conclude.
@@
  • If $\beta=\omega^{\beta_{0}}+\ldots+\omega^{\beta_{n}}$ with $n>0$ and $\beta_{0}\geq\ldots\geq\beta_{n}$, let $\operatorname{Sub}(\beta)=\{\beta\}+\operatorname{Sub}(\omega^{\beta_{0}})+\ldots+\operatorname{Sub}(\omega^{\beta_{n}})$.
Notice that if $\alpha\in\operatorname{Sub}(\beta)$ and $\beta\in\operatorname{Sub}(\gamma)$ then $\alpha\in\operatorname{Sub}(\gamma)$. Moreover, the multiplicity of $\alpha$ in the multiset $\operatorname{Sub}(\gamma)$ is at least that of $\alpha$ in $\operatorname{Sub}(\beta)$.
@@
$\bar{p}_{1}(\beta_{0},\beta_{1},\dots,\beta_{l})=(\overline{\beta_{0},\beta_{1}},\overline{\beta_{1},\beta_{2}},\ldots,\overline{\beta_{l},0}).$
@@
The proof is now by induction on $\tau$. Notice that $0\in S(\tau)$ for every $\tau$, which takes care of the case when $\tau=A(0)\leq A(1)$. So we can assume $A(0)>A(1)$.
@@
When $\nu=\gamma$ this shows that the coloring $\overline{c}$ defined above can actually be regarded as a coloring of the $(1+\gamma)$-size subsets of $s$. This is because in the coloring $\overline{c}$ of a set $u$ we only care about the peeling functions of index $\gamma$ or index $\nu\in S(T(\min u))$. By definition of the set $M$ (which contains $u$) each element of $u$ is larger than $T(\min u)||$ and by definition of $||\cdot||$, it is also larger than $|\nu|$ for each $\nu\in S(T(\min u))$. This means that if $u$ is $\gamma$-large, since $\gamma\geq\nu$, $u$ must also be $\nu$-large and so Lemma 6.13 yields that also $p_{\nu}$ only depends on the $(1+\nu)$-size initial segment of its input.
@@
By definition of $p_{\zeta}(t^{-0})$, we know that the Cantor normal forms of the ordinals $p_{\zeta-1}(t^{-0})$ and $p_{\zeta-1}(t^{-1})$ are equal up to a certain term $\omega^{\alpha_{n_{0}}}$, and then in $p_{\zeta-1}(t^{-0})$ we have the term $\omega^{p_{\zeta}(t^{-0})}$, while the tail $\tau_{1}$ of $p_{\zeta-1}(t^{-1})$ is strictly smaller than $\omega^{p_{\zeta}(t^{-0})}$. Since $p_{\zeta}(t^{-0})\leq p_{\zeta}(t^{-1})$, we must have that $p_{\zeta}(t^{-1})$ is an exponent of the Cantor normal form of $p_{\zeta-1}(t^{-1})$ that shows up before $n_{0}$. This implies that $p_{\zeta}(t^{-1})$ is one of the exponents of $p_{\zeta-1}(t^{-0})$ too. Since $\omega^{p_{\zeta}(t^{-0})}>\tau_{1}$, we must have that $\omega^{p_{\zeta}(t^{-1})}$ comes strictly before $\omega^{p_{\zeta}(t^{-0})}$ in the Cantor normal form of $p_{\zeta-1}(t^{-0})$. That is, even if they are equal as ordinals, they are different subterms of $^{p_{\zeta}(t^{-0})}$. Continuing like this we get that each $p_{\zeta}(t^{-i})$ is a different exponent in the Cantor normal form of $p_{\zeta-1}(t^{-0})$}. Finally, recall that $p_{\zeta-1}(t^{-0})$} is already a subterm of $T(t(0))$ by Lemma 6.5. Therefore, we get that that each $p_{\zeta}(t^{-i})$ for $i\leq|t_{\alpha}|$ originates from a distinct occurrence of an element of $\operatorname{Sub}(T(t(0)))$, as required.